These are Russian nesting dolls. The first doll is 5 inches, the second doll is 4. If the height of each consecutive doll is scaled down by the same ratio, Find the height of the third doll. ( hint the answer isn;t 3 inches)
Answer by nerdybill(7384)

These are Russian nesting dolls. The first doll is 5 inches, the second doll is 4. If the height of each consecutive doll is scaled down by the same ratio, Find the height of the third doll. ( hint the answer isn;t 3 inches)
.
Let x = ratio
then
5x = 4
x = 4/5
.
third doll then is:
4(4/5) = 16/5 = 3.2 inches